I am confused with what the question is asking please help.
what is the equation in standard form. points are (-3,2)and(1,2)

Answer by Nate(3500)

the equation for slope is the difference of the y value over the difference of the x values
m = (1st(y) - 2nd(y))/(1st(x) - 2nd(x)) = (2-2)/(1+3) = (0)/(4) = 0
m=0
the equation for the equation is:
y - (1st(y))=m(x - (1st(x)))
y - 2 = 0 (x + 3)
y - 2 = 0
y = 2